OpenStudy (anonymous):

Solve -3x2 + 30x - 90 = 0. solve with completing the square or quadratic formula. the answer should be a complex number i will reward a medal

OpenStudy (gorv):

multiply with -1 3x2-30x+90=0 divide by 3 x2-10x+30=0 x2-2*5*x+25+5=0 (x-5)^2=-5 x-5=(-5)^1/2 x=5+i*5^1/2

OpenStudy (gorv):

@marinesrock

OpenStudy (anonymous):

lets use quadratic formula: \[\frac{ -30 \pm \sqrt{30^{2}-4(-3)(-90)} }{ 2(-3) }\] reduces to: \[\frac{ -30 \pm \sqrt{-180} }{ -6 }\] take the negative out of square root with i and you get: x = 2.76i and 7.24i

OpenStudy (anonymous):

ok thanks